# Movable Antenna Enhanced Networked Full-Duplex Integrated Sensing and Communication System

### v2

@@ III-G Optimizing The TBS MAs’ Position
where $c_{4,5,m,k,n}$ is a constant term, $\delta_{1,m,k,n}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{18}_{m,k,n}[l]|$, $\mathbf{f}^{19}_{m,k,n}\triangleq\nabla h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})-\delta_{1,m,k,n}\mathbf{t}^{0}_{m,n,0}$, the elements of $\nabla h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})=\big{[}\frac{\partial h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})}{\partial x^{0}_{m,n,0}},\frac{\partial h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})}{\partial y^{0}_{m,n,0}}\big{]}^{T}$ are respectively formulated as
$\frac{\partial h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})}{\partial x^{0}_{m,n,0}}=-{\sum}_{l=1}^{L^{t,0}_{m,k}}\frac{2\pi}{\lambda}|\mathbf{f}^{18}_{m,k,n}[l]|\cos(\theta^{t,0}_{m,k,l})$ (65)
$\sin(\phi^{t,0}_{m,k,l})\sin\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{18}_{m,k,n}[l]\big{)},$
$\frac{\partial h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})}{\partial y^{0}_{m,n,0}}=-{\sum}_{l=1}^{L^{t,0}_{m,k}}\frac{2\pi}{\lambda}|\mathbf{f}^{18}_{m,k,n}[l]|\sin(\theta^{t,0}_{m,k,l})$ (66)
$\sin\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{18}_{m,k,n}[l]\big{)}.$
Similarly, the terms $\textrm{Re}\{({f}^{10}_{m,j,n})^{\ast}\mathbf{g}_{t,m,j}[n](\mathbf{t}^{0}_{m,n})\}$ and $\textrm{Re}\{(f^{13}_{i,m,n})^{\ast}\mathbf{g}_{t,m,i}[n](\mathbf{t}^{0}_{m,n})\}$ can be respectively convexified by leveraging (63) as follows
$\textrm{Re}\{({f}^{10}_{m,j,n})^{\ast}\mathbf{g}_{t,m,j}[n](\mathbf{t}^{0}_{m,n})\}=|{f}^{10}_{m,j,n}||\mathbf{g}_{t,m,j}[n]|$ (67)
$\cos\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{4}_{m,j}-\angle{f}^{10}_{m,j,n}\big{)}\triangleq g_{t1,m,j,n}(\mathbf{t}^{0}_{m,n})$
$\leq\frac{\delta_{2,m,j,n}}{2}\|\mathbf{t}^{0}_{m,n}\|^{2}_{2}+(\mathbf{f}^{20}_{m,j,n})^{T}\mathbf{t}^{0}_{m,n}+c_{4,6,m,j,n},$
$\textrm{Re}\{(f^{13}_{i,m,n})^{\ast}\mathbf{g}_{t,m,i}[n](\mathbf{t}^{0}_{m,n})\}=|f^{13}_{i,m,n}||\mathbf{g}_{t,m,i}[n]|$ (68)
$\cos\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{4}_{m,i}-\angle f^{13}_{i,m,n}\big{)}\triangleq g_{t2,i,m,n}(\mathbf{t}^{0}_{m,n})$
$\leq\frac{\delta_{4,i,m,n}}{2}\|\mathbf{t}^{0}_{m,n}\|^{2}_{2}+(\mathbf{f}^{22}_{i,m,n})^{T}\mathbf{t}^{0}_{m,n}+c_{4,7,i,m,n},$
with $c_{4,6,m,j,n}$ and $c_{4,7,i,m,n}$ being constants, $\delta_{2,m,j,n}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|{f}^{10}_{m,j,n}||\mathbf{g}_{t,m,j}[n]|$, $\delta_{4,i,m,n}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|f^{13}_{i,m,n}||\mathbf{g}_{t,m,i}[n]|$, $\mathbf{f}^{20}_{m,j,n}\triangleq\nabla g_{t1,m,j,n}(\mathbf{t}^{0}_{m,n})-\delta_{2,m,j,n}\mathbf{t}^{0}_{m,n,0}$, $\mathbf{f}^{22}_{i,m,n}\triangleq\nabla g_{t2,i,m,n}(\mathbf{t}^{0}_{m,n})-\delta_{4,i,m,n}\mathbf{t}^{0}_{m,n,0}$, $\nabla g_{t1,m,j,n}(\mathbf{t}^{0}_{m,n})$ and $\nabla g_{t1,m,j,n}(\mathbf{t}^{0}_{m,n})$ are the gradient vectors, and the details of their derivations are similar to (65) and (66) and are omitted to avoid repetition.
And then, we proceed to cope with non-convex term $\textrm{Re}\{(f^{16}_{i,m,n})^{\ast}\mathbf{g}_{t,m,0}[n](\mathbf{t}^{0}_{m,n})\}$. Unlike the upper-bounding procedures presented in (64) to (68), this time we require a convex lower bound for $\textrm{Re}\{(f^{16}_{m,n})^{\ast}\mathbf{g}_{t,m,0}[n](\mathbf{t}^{0}_{m,n})\}$ by applying (62) in Lemma 1, which is given as follows
$\textrm{Re}\{(f^{16}_{m,n})^{\ast}\mathbf{g}_{t,m,0}[n](\mathbf{t}^{0}_{m,n})\}=|f^{16}_{m,n}||\mathbf{g}_{t,m,0}[n]|$ (69)
$\cos\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{4}_{m,0}-\angle f^{16}_{m,n}\big{)}\triangleq g_{t3,m,n}(\mathbf{t}^{0}_{m,n})$
$\geq-\frac{\delta_{6,m,n}}{2}\|\mathbf{t}^{0}_{m,n}\|^{2}_{2}+(\mathbf{f}^{24}_{m,n})^{T}\mathbf{t}^{0}_{m,n}+c_{4,8,m,n},$
@@
where $c_{4,8,m,n}$ is constant, $\delta_{6,m,n}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|f^{16}_{m,n}||\mathbf{g}_{t,m,0}[n]|$, $\mathbf{f}^{24}_{m,n}\triangleq\nabla g_{t3,m,n}(\mathbf{t}^{0}_{m,n})+\delta_{6,m,n}\mathbf{t}^{0}_{m,n,0}$, $\nabla g_{t3,m,n}(\mathbf{t}^{0}_{m,n})$ is the the gradient vector, and the details of its derivation (refer to (65) and (66)) are omitted to avoid repetition.
@@
$\|\mathbf{t}^{0}_{m,n}-\mathbf{t}^{0}_{m,i}\|_{2}\geq\frac{(\mathbf{t}^{0}_{m,n,0}-\mathbf{t}^{0}_{m,i})^{T}}{\|\mathbf{t}^{0}_{m,n,0}-\mathbf{t}^{0}_{m,i}\|_{2}}(\mathbf{t}^{0}_{m,n}-\mathbf{t}^{0}_{m,i}).$ (70)
Based on the above MM transformation, the update of $\mathbf{t}^{0}_{m,n}$ can be performed by solving a convex surrogate of the original challenging problem (P9), as follows
[…]
s.t. $\delta_{7,m,n}\|\mathbf{t}^{0}_{m,n}\|^{2}_{2}\!+\!(\mathbf{f}^{25}_{m,n})^{T}\mathbf{t}^{0}_{m,n}\!+\!c_{4,10,m,n}\leq 0,$ (71b)
$\frac{(\mathbf{t}^{0}_{m,n,0}\!\!-\!\mathbf{t}^{0}_{m,i})^{T}}{\|\mathbf{t}^{0}_{m,n,0}\!\!-\!\mathbf{t}^{0}_{m,i}\|_{2}}\!(\mathbf{t}^{0}_{m,n}\!\!\!-\!\!\mathbf{t}^{0}_{m,i})\!\!\geq\!\!D_{t},\!\forall\!n,\!i\!\!\in\!\!\mathcal{N}_{t},n\!\!\neq\!\!i,$ (71c)
$\mathbf{t}^{0}_{m,n}\in\mathcal{C},$ (71d)
with $c_{4,9,m,n}$ and $c_{4,10,m,n}$ being constant terms, $\delta_{3,m,n}\triangleq\sum_{k=1}^{K_{d}}\frac{\delta_{1,m,k,n}}{2}+\sum_{j=0}^{K_{t}}\frac{\delta_{2,m,j,n}}{2}$, $\delta_{7,m,n}\triangleq\sum_{j=1}^{K_{t}}\frac{\delta_{4,j,m,n}}{2}+\frac{\delta_{6,m,n}}{2}$, $\mathbf{f}^{21}_{m,n}\triangleq\sum_{k=1}^{K_{d}}\mathbf{f}^{19}_{m,k,n}+\sum_{j=0}^{K_{t}}\mathbf{f}^{20}_{m,j,n}$ and $\mathbf{f}^{25}_{m,n}\triangleq\sum_{j=1}^{K_{t}}\mathbf{f}^{22}_{j,m,n}-\mathbf{f}^{24}_{m,n}$. Clearly, problem (P10) is convex and can be numerically solved.

### v3

@@ III-G Optimizing The TBS MAs’ Position
where $c_{4,6,k}$ is a constant term, $\delta_{1,k}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{7}_{k}[l]|$, $\mathbf{f}^{8}_{k}\triangleq\nabla h^{1}_{k}(\mathbf{t}^{0}_{0})-\delta_{1,k}\mathbf{t}^{0}_{0}$, and the elements of the gradient vector $\nabla h^{1}_{k}(\mathbf{t}^{0}_{0})=\big[\frac{\partial h^{1}_{k}(\mathbf{t}^{0}_{0})}{\partial x^{0}_{0}},\frac{\partial h^{1}_{k}(\mathbf{t}^{0}_{0})}{\partial y^{0}_{0}}\big]^{T}$ are respectively given as  [^6] In the following, when applying the equations (64) and/or (65), the derivation details of the gradient vector can refer to (67) and (68).
$\frac{\partial h^{1}_{k}(\mathbf{t}^{0}_{0})}{\partial x^{0}_{0}}=-{\sum}_{l=1}^{L^{t,0}_{m,k}}\frac{2\pi}{\lambda}|\mathbf{f}^{7}_{k}[l]|\cos(\theta^{t,0}_{m,k,l})$ (67)
$\sin(\phi^{t,0}_{m,k,l})\sin\big(\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{0})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{7}_{k}[l]\big),$
$\frac{\partial h^{1}_{k}(\mathbf{t}^{0}_{0})}{\partial y^{0}_{0}}=-{\sum}_{l=1}^{L^{t,0}_{m,k}}\frac{2\pi}{\lambda}|\mathbf{f}^{7}_{k}[l]|\sin(\theta^{t,0}_{m,k,l})$ (68)
$\sin\big(\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{0})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{7}_{k}[l]\big).$
Furthermore, the non-convex term $\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}\}$ can be lower-bounded by the inequality (64), which is given as
$\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}\}\triangleq h^{2}_{k}(\mathbf{t}^{0})$ (69)
$={\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{2}_{k}[l]|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{0})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{2}_{k}[l]\big)$
$\geq h^{2}_{k}(\mathbf{t}^{0}_{0})+\nabla^{T}h^{2}_{k}(\mathbf{t}^{0}_{0})(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})-0.5{\bar{\delta}_{1,k}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})^{T}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})$
$=-0.5{\bar{\delta}_{1,k}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{{f}}^{9}_{k})^{T}\mathbf{t}^{0}+{c}_{4,7,k},$
@@
where ${c}_{4,7,k}$ is a constant term, $\bar{\delta}_{1,k}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{2}_{k}[l]|$, $\mathbf{{f}}^{9}_{k}\triangleq\nabla h^{2}_{k}(\mathbf{t}^{0}_{0})+\bar{\delta}_{1,k}\mathbf{t}^{0}_{0}$, $\nabla h^{2}_{k}(\mathbf{t}^{0}_{0})$ is the gradient vector.
Next, the term $\textrm{Re}\{({f}^{3}_{j})^{\ast}{g}_{t,j}(\mathbf{t}^{0})\}$ be convexified by leveraging (65) as follows
$\textrm{Re}\{({f}^{10}_{j})^{\ast}{g}_{t,j}(\mathbf{t}^{0})\}\triangleq g^{1}_{j}(\mathbf{t}^{0})$ (70)
$=|{f}^{3}_{j}||{g}_{t,j}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{0})^{T}\mathbf{a}^{4}_{m,j}-\angle{f}^{3}_{j}\big)$
$\leq g^{1}_{j}(\mathbf{t}^{0}_{0})+\nabla^{T}g^{1}_{j}(\mathbf{t}^{0}_{0})(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})+0.5{\delta_{2,j}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})^{T}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})$
$=0.5{\delta_{2,j}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{10}_{j})^{T}\mathbf{t}^{0}+c_{4,8,j},$
with $c_{4,8,j}$ being constant, $\delta_{2,j}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|{f}^{3}_{j}||\mathbf{g}_{t,j}|$, $\mathbf{f}^{10}_{j}\triangleq\nabla g^{1}_{j}(\mathbf{t}^{0}_{0})-\delta_{2,j}\mathbf{t}^{0}_{0}$, $\nabla g^{1}_{j}(\mathbf{t}^{0}_{0})$ is the gradient vector over $\mathbf{t}^{0}_{0}$.
Based on the above transformation, the upper-bound of the objective function (61a) is given as
${\sum}_{k=1}^{K_{d}}\big((\mathbf{h}^{0}_{k}(\mathbf{t}^{0}))^{H}\mathbf{{F}}_{1,k}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})+\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}$ (71)
$-\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}\big)+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{({f}^{3}_{j})^{\ast}{g}_{t,j}(\mathbf{t}^{0})\}+\hat{c}_{4,2}$
$\leq{\sum}_{k=1}^{K_{d}}\big(0.5{\delta_{1,k}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{8}_{k})^{T}\mathbf{t}^{0}+c_{4,6,k}$
$-(-{0.5\bar{\delta}_{1,k}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{{f}}^{9}_{k})^{T}\mathbf{t}^{0}+{c}_{4,7,k})\big)$
$+{\sum}_{j=0}^{K_{t}}({0.5\delta_{2,j}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{10}_{j})^{T}\mathbf{t}^{0}+c_{4,8,j})+{c}_{4,4}$
$={\delta_{3}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{11})^{T}\mathbf{t}^{0}+c_{4,9}$
where $c_{4,9}$ is a constant term, $\delta_{3}\triangleq\sum_{k=1}^{K_{d}}({\delta_{1,k}}/{2}+{\bar{\delta}_{1,k}}/{2})+\sum_{j=0}^{K_{t}}{\delta_{2,j}}/{2}$ and $\mathbf{f}^{11}\triangleq\sum_{k=1}^{K_{d}}(\mathbf{f}^{8}_{k}-\mathbf{{f}}^{9}_{k})+\sum_{j=0}^{K_{t}}\mathbf{f}^{10}_{j}$.
In the following, the term $\textrm{Re}\{(f^{13}_{i})^{\ast}{g}_{t,i}(\mathbf{t}^{0})\}$ of the sensing SINR constraint (61b) can be convexified via (65) as follows
$\textrm{Re}\{(f^{4}_{i})^{\ast}\mathbf{g}_{t,i}(\mathbf{t}^{0})\}\triangleq g^{2}_{i}(\mathbf{t}^{0})$ (72)
$=|f^{4}_{i}||\mathbf{g}_{t,i}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{0})^{T}\mathbf{a}^{4}_{m,i}-\angle f^{4}_{i}\big)$
$\leq g^{2}_{i}(\mathbf{t}^{0}_{0})+\nabla^{T}g^{2}_{i}(\mathbf{t}^{0}_{0})(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})+0.5{\delta_{4,i}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})^{T}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})$
$={0.5\delta_{4,i}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{12}_{i})^{T}\mathbf{t}^{0}+c_{4,10,i},$
with $c_{4,10,i}$ being constant, $\delta_{4,i}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|f^{4}_{i}||\mathbf{g}_{t,i}|$, $\mathbf{f}^{12}_{i}\triangleq\nabla g^{2}_{i}(\mathbf{t}^{0}_{0})-\delta_{4,i}\mathbf{t}^{0}_{0}$, $\nabla g^{2}_{i}(\mathbf{t}^{0}_{0})$ is the gradient vector.
And then, we proceed to cope with the non-convex term $\textrm{Re}\{(f^{5})^{\ast}{g}_{t,0}(\mathbf{t}^{0})\}$ of the sensing SINR constraint (61b). This time, we require a convex lower bound for $\textrm{Re}\{(f^{5})^{\ast}{g}_{t,0}(\mathbf{t}^{0})\}$ by applying (64) in Lemma 1, which is given as follows
$\textrm{Re}\{(f^{5})^{\ast}{g}_{t,0}(\mathbf{t}^{0})\}\triangleq g^{3}(\mathbf{t}^{0})$ (73)
$=|f^{5}||\mathbf{g}_{t,0}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{0})^{T}\mathbf{a}^{4}_{m,0}-\angle f^{5}\big)$
$\geq g^{3}(\mathbf{t}^{0}_{0})+\nabla^{T}g^{3}(\mathbf{t}^{0}_{0})(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})-{0.5\delta_{6}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})^{T}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})$
$=-{0.5\delta_{6}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{13})^{T}\mathbf{t}^{0}+c_{4,11},$
where $c_{4,11}$ is constant, $\delta_{6}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|f^{5}||\mathbf{g}_{t,0}|$, $\mathbf{f}^{13}\triangleq\nabla g^{3}(\mathbf{t}^{0}_{0})+\delta_{6}\mathbf{t}^{0}_{0}$, $\nabla g^{3}(\mathbf{t}^{0}_{0})$ is the gradient vector.
Therefore, a convex upper-bound of the SINR constraint (61b) can be written as
${\sum}_{i=1}^{K_{t}}\textrm{Re}\{(f^{4}_{i})^{\ast}{g}_{t,i}(\mathbf{t}^{0})\}-\textrm{Re}\{(f^{5})^{\ast}{g}_{t,0}(\mathbf{t}^{0})\}+c_{4,3}$ (74)
$\leq{\sum}_{i=1}^{K_{t}}({\delta_{4,i}}/{2}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{12}_{i})^{T}\mathbf{t}^{0}+c_{4,10,i})$
$-(-{\delta_{6}}/{2}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{13})^{T}\mathbf{t}^{0}+c_{4,11})+c_{4,3}$
$=\delta_{7}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{14})^{T}\mathbf{t}^{0}+c_{4,12},$
where $c_{4,12}$ is constant term, $\delta_{7}\triangleq\sum_{j=1}^{K_{t}}{\delta_{4,j}}/{2}+{\delta_{6}}/{2}$ and $\mathbf{f}^{14}\triangleq\sum_{j=1}^{K_{t}}\mathbf{f}^{12}_{j}-\mathbf{f}^{13}$.
@@
$\|\mathbf{t}^{0}-\mathbf{t}^{0}_{m,i}\|_{2}\geq\frac{(\mathbf{t}^{0}_{0}-\mathbf{t}^{0}_{m,i})^{T}}{\|\mathbf{t}^{0}_{0}-\mathbf{t}^{0}_{m,i}\|_{2}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{m,i}).$ (75)
Based on the above MM transformation, the update of $\mathbf{t}^{0}$ can be performed by solving a convex surrogate of the original challenging problem (P9), as follows
$\textrm{(P10)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{t}^{0}}\delta_{3}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{11})^{T}\mathbf{t}^{0}_{m,n}+c_{4,9}$ (76a)
s.t. $\delta_{7}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{14})^{T}\mathbf{t}^{0}+c_{4,12}\leq 0,$ (76b)
$\frac{(\mathbf{t}^{0}_{0}-\mathbf{t}^{0}_{m,i})^{T}}{\|\mathbf{t}^{0}_{0}\!\!-\!\mathbf{t}^{0}_{m,i}\|_{2}}\!(\mathbf{t}^{0}\!-\!\mathbf{t}^{0}_{m,i})\geq D_{t},\forall i\in\mathcal{N}_{t},n\neq i,$ (76c)
$\mathbf{t}^{0}\in\mathcal{C},$ (76d)
Clearly, problem (P10) is convex and can be numerically solved.